# A Scalable Bayesian Spatiotemporal Model for Water Level Predictions using a Nearest Neighbor Gaussian Process Approach

### v1

@@ 3 Methodology
The proposed model is inspired by a study on Australian river water temperatures by Santos-Fernandez et al., (2022), which was selected due to its suitability for addressing the challenges posed by long-term spatiotemporal data (42 spatial locations measured over 87 days), a time dimension similar to the current study. The authors proposed a Bayesian spatiotemporal model for predicting water temperature where the temporal correlation is captured through an autoregressive parameter and the spatial dependence along the river network is modeled using a Gaussian Process (GP) that incorporates stream distance. In contrast, our study deals with a significantly larger dataset comprising 301 spatial locations compared to their 42. This substantial increase in spatial points makes the use of a conventional GP computationally impractical due to the inversion of a large spatial covariance matrix. In order to address the computational limitation, we employed the Nearest Neighbor Gaussian Process (NNGP), a scalable alternative that effectively approximates the GP while preserving spatial dependencies. First, we present the concept of NNGP in Section 3.1. Subsequently, we define the proposed spatiotemporal model in Section 3.2.
@@ 3.1 Nearest Neighbor Gaussian Process
Let $\boldsymbol{s}=\{\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{n}\}$ be the set of spatial locations in the spatial domain $\mathcal{S}\subset\mathbb{R}^{d}$ and consider a pure spatial process following a zero-centered GP with covariance matrix $\boldsymbol{\Sigma}=\Sigma(\cdot,\cdot)$, denoted as $\boldsymbol{w}\sim GP(\boldsymbol{0},\boldsymbol{\Sigma})$. The process can be expressed as the product of conditional normal distributions
@@
In a setting with a substantial number of spatial locations, some elements within the conditional set provide a superfluous contribution, in such a manner that their exclusion results in minimal loss of information. The underlying idea of the NNGP is to substitute each extensive conditional set by a subset (smaller set) containing at most $m$ nearest neighbors (Vecchia,, 1988). Formally, let $N(\boldsymbol{s}_{i})$ be the subset of $m$ nearest neighbors of location $\boldsymbol{s}_{i}$, that is, $N(\boldsymbol{s}_{i})\subset\{\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{i-1}\}$. We have an approximation for Equation 1, known as the Nearest Neighbor Gaussian Process (NNGP) denoted by $\widetilde{\boldsymbol{w}}\sim NNGP(\boldsymbol{0},\widetilde{\boldsymbol{\Sigma}})$
@@
As demonstrated by Datta et al., 2016a the stated approximation is a valid multivariate normal distribution and also is a valid spatial process over $\mathcal{S}$. Hence the predictions are close to the GP model and the NNGP kriging recovers the true spatial process. Defining the neighbor set $N(\boldsymbol{s}_{i})$ requires imposing an order on locations, such that any given order generates a valid joint distribution. We followed the approach as suggested by Datta et al., 2016a , ordering according to one of the spatial coordinates, providing a convenient and effective solution. Indeed, we ordered the spatial locations by the latitude coordinate.
@@
The presented NNGP was formulated as a zero-centered distribution but it is trivial to shift the mean to accommodate a non-zero centered process. Furthermore, the process can be simply extended to a space-time setting, with the incorporation of a separable spatiotemporal covariance model, as detailed in Section 3.2.
@@ 3.2 The spatiotemporal model
Let $\boldsymbol{y}_{t}$ be the vector of water level measurements for all locations at time $t=1,\dots,T$ and the stacked vector $[\boldsymbol{y}_{1},\boldsymbol{y}_{2},\dots,\boldsymbol{y}_{T}]$ representing the multivariate distribution for all locations and time points. We propose a model based on the conditional distributions
@@
where $\boldsymbol{X}_{t}$ is the design matrix of predictors at time $t$, $\boldsymbol{\theta}$ is the vector of parameters and $\boldsymbol{y}_{1}$ is the distribution at time $t=1$. For the following time points the distribution is given by
@@
A full Bayesian inference approach was employed to naturally handle missing values and provide thorough quantification of uncertainty. The model was implemented in Stan using Markov Chain Monte Carlo (MCMC) sampling via No-U-Turn Sampler (NUTS). In order to facilitate the interface with Stan, the cmdstanr package in R was used (Gabry et al.,, 2024). The implementation efficiently completed 500 iterations across 4 chains in approximately 17 minutes. The code used for model implementation is available in this Github repository. All scripts and instructions necessary to reproduce the results are provided.
@@ 3.3.1 Unobserved (new) location at a past time point
$y_{t}(\boldsymbol{s}_{new})\mid y_{t-1}(\boldsymbol{s}_{new}),\boldsymbol{y}_{t},\boldsymbol{y}_{t-1}\sim\mathcal{N}(\alpha,\gamma)$
$\alpha=\mu_{t}(\boldsymbol{s}_{new})+\Sigma(\boldsymbol{s}_{new},N(\boldsymbol{s}_{new}))(\sigma^{2})^{-1}(\boldsymbol{y}_{t}(N(\boldsymbol{s}_{new}))-\boldsymbol{\mu}_{t}(N(\boldsymbol{s}_{new})))$
$\gamma=\sigma^{2}-\Sigma(\boldsymbol{s}_{new},N(\boldsymbol{s}_{new}))\boldsymbol{\Sigma}_{N}^{-1}\Sigma(\boldsymbol{s}_{new},N(\boldsymbol{s}_{new}))^{T}.$
@@ 4 Application
In terms of performance, the goal is to make predictions at locations where there is no data available, as well as to forecast future observations at sites with existing data. In order to evaluate the model performance in both scenarios, an out-of-sample (OOS) evaluation was conducted, splitting the data into two subsets: training and validation. Specifically, the validation dataset consisted of two components: (1) a random selection of stations that included all available time points, and (2) the reservation of the last three days of observations for the remaining stations. The illustrative diagram in Figure 7 represents the data splitting strategy for a total of 20 stations.
The Section 4.1 first presents the data generation procedure with missing values and the application of the mentioned splitting setup. Subsequently, it outlines the results obtained by the proposed model. The Section 4.2 presents the validation results of the proposed model for the water level dataset, where 20 stations were randomly selected for testing. This left a training set consisting of 281 stations, each measured over a period of 85 days, resulting in a total of 27,090 observations.
Figure 7: Illustrative diagram of training and validation dataset for out-of-sample evaluation. Each tile represents to a daily observation and coloured according to the data subset.
@@ 4.1 Synthetic dataset
We simulated a spatiotemporal dataset composed of 100 spatial locations within a unit square domain, each measured for 20 days, in a total of $100\times 20=2,000$ observations. First we simulated the spatial locations in a unit square. Following that, the spatial covariance matrix was computed using the exponential function with parameters $\sigma^{2}=1$ and $l=0.3$. The response variable was then simulated sequentially from $t=1$ onward, using Equation 6. The mean function (Equation 7) assumed regression coefficients $\boldsymbol{\beta}=[1,3]$ and $\phi=0.9$. The parameter values were selected in a manner that ensured the temporal correlation resulted in a long-lasting effect while simultaneously preventing the spatial parameters from being overly smoothed or localised. Finally, the space-time process was fully specified adding the error variance ($\tau^{2}=0.1$) into the spatial covariance. We can summarise the generation of the synthetic dataset as
@@
Figure 8: Predictive performance of the proposed model for the synthetic dataset considering prediction 1-day ahead under two different scenarios (unseen and observed locations). The dots represents the posterior mean while the vertical bars represent the 95% prediction interval.
@@ 4.2 Water level dataset
Table 2 shows the posterior estimates for the water level dataset. The large value of $\phi$ (0.89) highlights the high temporal dependence between the water level in consecutive days as our exploratory analysis previously indicated in Figure 4. In other words, the water level in the future days is similar to the previous day if there is no large changes in precipitation. The posterior mean of the length-scale parameter is $l=24.53$ kilometers, which is a considerable range for the spatial decay. This range indicates that locations within approximately 25 kilometers of each other are correlated, which is consistent with the physical continuity of river networks and the gradual change in water levels over short distances. Furthermore, the proportion of variance that is captured by the spatial latent field - $\sigma^{2}/(\sigma^{2}+\tau^{2})\approx 0.76$ emphasising the importance of including the spatial component.
@@
$\beta_{0}$ 0.01 0.01 (-0.02, 0.04)
$\beta_{1}$ 0.05 0.05 (0.05, 0.99)
$\phi$ 0.89 0.89 (0.88, 0.89)
$\sigma^{2}$ 0.01 0.01 (0.01, 0.01)
$\tau^{2}$ 0.003 0.003 (0.003, 0.004)
$l$ 24.53 24.55 (23.22, 25.74)
After obtaining samples from the posterior distribution, they are used to predict the water levels at locations without available data and at future times. Figure 10 illustrates the comparison of the predicted and the true latent water level when predicting at one, two or three days in advance. The findings demonstrated that the model accurately predicts the true water level when predicting one and two days ahead. On the other hand, the prediction for three days ahead demonstrated inferior performance, with an underestimation of high water level records exceeding 0.4 meters. As expected, the prediction for an observed location on a future day exhibits superior accuracy and more narrow prediction intervals in comparison to sites without data (Table 3).
@@
Figure 10: Predictive performance of the proposed model for the water level dataset considering predictions 1, 2 and 3-days ahead under two different scenarios (unseen and observed locations). The dots represents the posterior mean and while the vertical bars represent the 95% prediction interval.
@@
Table 3: RMSE of OOS evaluation for prediction in 1, 2, 3-days ahead according to data availability.
Date Observed locations Unseen locations
02/04/2022 (1-day ahead) 0.039 0.141
03/04/2022 (2-day ahead) 0.081 0.155
04/04/2022 (3-day ahead) 0.129 0.160
@@
Figure 11 depicts the predicted water level (after backtransformation) for 20 unobserved locations during the process of model fitting. At each observed time point, an interpolation was performed to estimate water levels at these unobserved locations. For dates extending beyond the training dataset (02/04/2022 to 06/04/2022), the spatial latent field was first computed for the observed locations and subsequently interpolated in order to predict the water level at the unobserved locations, as outlined in Section 3.3. The limited information beyond 02/04/2022 is reflected in wider prediction intervals, indicating greater uncertainty in predictions.
In general, the model demonstrated the capacity to predict the water level for unobserved locations, including on a future day. However, for stations (18002, 18019, 25001 and 23030) the model struggles to accurately predict peak levels, though capturing well the general variation. Additionally, in some cases, the model overestimates the water level fluctuations (stations 14121 and 22009). These inconsistencies indicates that while the model exhibits satisfactory performance under typical circumstances, it may require additional refinement to handle site-specific characteristics.
@@
Figure 11: Predicted water level and its true values for 20 unobserved stations. The black dots represent the real values, while the blue line represents the posterior mean with the shaded area indicating the 95% prediction interval.
@@ 5 Model Comparison
Despite the model’s novelty in the hydrology field, we aim to compare its prediction performance and uncertainty quantification with other methodologies in the literature within two different setups: (1) comparison with a competitor model for large spatiotemporal datasets and (2) comparison with models focused on a single basin, which focus on river systems draining into a single large river, which is the approach commonly used by hydrologists. This dual comparison allows us to assess the model’s scalability and effectiveness across diverse hydrological contexts.
@@ 5.2 Comparison with a single-basin model
In this section, we investigate the advantages of the proposed model by comparing it to models that focus on a single basin, a common approach in hydrology (Varouchakis et al.,, 2022; Varouchakis and Hristopulos,, 2019; Kazemi et al.,, 2021; Ruybal et al.,, 2019). Single-basin models are convenient because they alleviate the computational bottleneck of inverting large covariance matrices while preserving the unique characteristics of the basin. However, they may limit the model’s ability to leverage information from multiple basins, which could enhance its performance. Our comparison aims to highlight the trade-off between computational feasibility and model robustness when scaling to larger, multi-basin datasets.
@@
5.2.1 GP 1: Autoregressive process into the dependent variable
Spatiotemporal GP model that the mean at current time depends on the response in the previous time.
$p(\boldsymbol{y}_{t}$ $\mid\boldsymbol{y}_{t-1},\boldsymbol{X}_{t},\boldsymbol{X}_{t-1},\boldsymbol{\theta})\sim\mathcal{N}(\boldsymbol{y}_{t};\boldsymbol{\mu}_{t},\boldsymbol{\Sigma}+\tau^{2}\boldsymbol{I})$ (11)
$\boldsymbol{\mu}_{t}$ $=\boldsymbol{X}_{t}\boldsymbol{\beta}+\phi(\boldsymbol{y}_{t-1}-\boldsymbol{X}_{t-1}\boldsymbol{\beta}).$
@@ 5.2.1 GP 1: Autoregressive process into the dependent variable
Spatiotemporal GP model that the temporal dependence is incorporated in the spatial process.
$p(\boldsymbol{y}_{t}$ $\mid\boldsymbol{X}_{t},\boldsymbol{z}_{t},\boldsymbol{\theta})\sim\mathcal{N}(\boldsymbol{y}_{t};\boldsymbol{X}_{t}\boldsymbol{\beta}+\boldsymbol{w}_{t},\tau^{2}\boldsymbol{I})$ (12)
$\boldsymbol{w}_{t}$ $=\rho\boldsymbol{w}_{t-1}+\boldsymbol{\xi}_{t},\quad\boldsymbol{\xi}_{t}\sim GP(\boldsymbol{0},\boldsymbol{\Sigma}).$

### v2

@@ 3 Methodology
The proposed model draws inspiration by a study on Australian river water temperatures conducted by Santos-Fernandez et al., (2022). This work was selected for its suitability in addressing the challenges posed by long-term spatiotemporal data - specifically, 42 spatial locations measured over 87 days. Notably, the time range closely resembles that of the present work. The authors proposed a Bayesian spatiotemporal model for predicting water temperature which the temporal autocorrelation is captured through an autoregressive parameter and the spatial dependence along the river network is modelled through a Gaussian Process (GP) depending on Euclidean and stream distance. In contrast, our study deals with a significantly larger dataset comprising 301 spatial locations compared to their 42. This substantial increase in spatial points makes the use of a conventional GP computationally impractical due to the inversion of a large spatial covariance matrix. In order to address the computational limitation, we employed the Nearest Neighbor Gaussian Process (NNGP), a scalable alternative that effectively approximates the GP while preserving spatial dependencies. First, we formalise the NNGP approach in Section 3.1. Subsequently, we present the proposed spatiotemporal model in Section 3.2.
@@ 3.1 Nearest Neighbor Gaussian Process
Let $\boldsymbol{s}=\{\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{n}\}$ be the set of spatial locations in the spatial domain $\mathcal{S}\subset\mathbb{R}^{d}$ and consider a pure spatial process following a zero-centred GP with covariance matrix $\boldsymbol{\Sigma}=\Sigma(\cdot,\cdot)$, denoted as $\boldsymbol{w}\sim GP(\boldsymbol{0},\boldsymbol{\Sigma})$. The process can be expressed as the product of conditional normal distributions
@@
In a setting with a substantial number of spatial locations, a certain elements in the conditional set offer minimal contribution, in such a manner that their exclusion results in negligible loss of information. The underlying idea of the NNGP is to substitute each extensive conditional set by a subset (smaller set) containing at most $m$-nearest neighbors (Vecchia,, 1988). Formally, let $N(\boldsymbol{s}_{i})$ be the subset of at most $m$-nearest neighbors with respect to location $\boldsymbol{s}_{i}$, that is, $N(\boldsymbol{s}_{i})\subset\{\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{i-1}\}$. We have an approximation for Equation 1, known as the Nearest Neighbor Gaussian Process (NNGP) denoted by $\widetilde{\boldsymbol{w}}\sim NNGP(\boldsymbol{0},\widetilde{\boldsymbol{\Sigma}})$
@@
As demonstrated by Datta et al., 2016a , the stated approximation is a valid multivariate normal distribution and also is a valid spatial process over $\mathcal{S}$. Hence, the predictions are close to the GP model and the NNGP kriging recovers the true spatial process. Defining the neighbor set $N(\boldsymbol{s}_{i})$ requires imposing an order on locations, such that any given order generates a valid joint distribution. Here, we ordered the locations based on the latitude coordinate, as suggested by Datta et al., 2016a .
@@
The presented NNGP was formulated as a zero-centred distribution but it is trivial to shift the mean to accommodate a non-zero centred process. Furthermore, the process can be simply extended to a space-time setting, with the incorporation of a separable spatiotemporal covariance structure, as detailed in Section 3.2.
@@ 3.2 The spatiotemporal model
Let $\boldsymbol{y}_{t}$ be the vector of observed water level for all locations at time $t=1,\dots,T$. Therefore, the stacked vector $[\boldsymbol{y}_{1},\boldsymbol{y}_{2},\dots,\boldsymbol{y}_{T}]$ represents the multivariate distribution for all locations and time points. We propose a model based on the product of conditional distributions
@@
where $\boldsymbol{X}_{t}$ is the design matrix of predictors at time $t$, $\boldsymbol{\theta}$ is the vector of parameters and $\boldsymbol{y}_{1}$ is the distribution at time $t=1$. For the following time points, the distribution is given by
@@
A full Bayesian inference approach was employed to naturally handle missing values and provide thorough quantification of uncertainty. The model was implemented in Stan using Markov Chain Monte Carlo (MCMC) sampling via No-U-Turn Sampler (NUTS). In order to facilitate the interface with Stan, the cmdstanr package in R was used (Gabry et al.,, 2024). The implementation efficiently completed 500 iterations across 4 chains in approximately 36 minutes. The code used for model implementation is available in this Github repository. All scripts and instructions necessary to reproduce the results are provided.
@@ 3.3.1 Unobserved (new) location at a past time point
$y_{t}(\boldsymbol{s}_{new})\mid y_{t-1}(\boldsymbol{s}_{new}),\boldsymbol{y}_{t},\boldsymbol{y}_{t-1}\sim\mathcal{N}(\mu_{new},\sigma^{2}_{new})$
$\mu_{new}=\mu_{t}(\boldsymbol{s}_{new})+\Sigma(\boldsymbol{s}_{new},N(\boldsymbol{s}_{new}))(\sigma^{2})^{-1}(\boldsymbol{y}_{t}(N(\boldsymbol{s}_{new}))-\boldsymbol{\mu}_{t}(N(\boldsymbol{s}_{new})))$
$\sigma^{2}_{new}=\sigma^{2}-\Sigma(\boldsymbol{s}_{new},N(\boldsymbol{s}_{new}))\boldsymbol{\Sigma}_{N}^{-1}\Sigma(\boldsymbol{s}_{new},N(\boldsymbol{s}_{new}))^{T}.$
@@ 4 Application
In terms of performance, the goal is to make predictions at locations where there is no data available, as well as to forecast future observations at sites with existing data. In order to evaluate the model performance in both scenarios, an out-of-sample (OOS) evaluation was conducted, splitting the data into two subsets: training and validation. Specifically, the validation dataset consisted of two components: (1) a random selection of stations that included all available time points, and (2) the reservation of the last three days of observations for the remaining stations. The illustrative diagram in Figure 7 represents the data splitting strategy for a hypothetical scenario with 20 stations.
Section 4.1 first presents the data generation procedure with missing values and the application of the mentioned splitting setup. Subsequently, it outlines the results obtained by the proposed model. Section 4.2 presents the validation results of the proposed model for the water level dataset, where 20 stations were randomly selected for testing. This left a training set consisting of 281 stations, each measured over a period of 85 days, resulting in a total of 27,090 observations.
Figure 7: Illustrative diagram of training and validation dataset for out-of-sample evaluation for a scenario with 20 monitoring stations. Three stations were randomly selected to contribute exclusively to the validation dataset, containing all their daily observations. Each tile represents a daily observation and is coloured-coded to indicate whether it belongs to training or validation set.
@@ 4.1 Synthetic dataset
We simulated a spatiotemporal dataset composed of 100 spatial locations within a unit square domain, each measured for 20 days, in a total of $100\times 20=2,000$ observations. First we simulated the spatial locations in a unit square. Following that, the spatial covariance matrix was computed using the exponential function with parameters $\sigma^{2}=1$ and $l=0.3$. The response variable was then simulated sequentially from $t=1$ onward, using Equation 6. The mean function (Equation 7) assumed regression coefficients $\boldsymbol{\beta}=[1,3]$ and autoregressive parameter $\phi=0.9$. The parameter values were selected in a manner that ensured the temporal correlation resulted in a long-lasting effect while simultaneously preventing the spatial parameters from being overly smoothed or localised. Finally, the space-time process was fully specified adding the error variance ($\tau^{2}=0.1$) into the spatial covariance. We can summarise the generation of the synthetic dataset as
@@
Figure 8: Predictive performance of the proposed model for the synthetic dataset considering prediction one-day ahead under two different scenarios (unseen and observed locations). The dots represents the posterior mean while the vertical bars represent the 95% prediction interval.
@@ 4.2 Water level dataset
Table 2 shows the posterior estimates for the water level dataset. The large value of the autoregressive parameter ($\phi=0.87$) highlights the strong temporal dependence between the water level in consecutive days. In other words, the water level in the future days is similar to the previous day if there is no large changes in precipitation. The posterior mean of the length-scale parameter is $l=34.14$ kilometers, which is a considerable range for the spatial decay. This range indicates that locations within approximately 35 kilometers of each other are correlated, which is consistent with the physical continuity of river networks and the gradual change in water levels over short distances. Furthermore, the proportion of variance that is captured by the spatial latent field - $\sigma^{2}/(\sigma^{2}+\tau^{2})\approx 0.76$ emphasising the importance of including the spatial component.
@@
$\beta_{0}$0.02 -0.02 (-0.06, 0.02)
$\beta_{1}$ 0.024 0.024 (0.021, 0.027)
$\phi$ 0.878 0.878 (0.874, 0.884)
$\sigma^{2}$ 0.014 0.014 (0.013, 0.014)
$\tau^{2}$ 0.0043 0.0043 (0.0041, 0.0044)
$l$ 34.14 34.14 (32.14, 36.35)
After obtaining samples from the posterior distribution, we used them to derive the posterior predictive distribution of water levels at locations without available observation and at future time points. Figure 10 presents the comparison between the predicted and true water level for forecasts ranging from one to five days ahead. The results demonstrate that the model performs well for short-term predictions, particularly for one and two-day in advance, as the posterior mean is close to the true values with narrow predictive intervals. However, the predictive performance decreases with longer forecast horizons, with the three-day-ahead prediction exhibiting noticeable overestimation and great uncertainty, as reflected by the wider prediction intervals. This increase in uncertainty is a expected result, as longer-term forecasts inherently carry more variability. Similarly, the prediction for an observed location on a future day exhibits superior accuracy and more narrow prediction intervals in comparison to sites without data (see Table 3).
@@
Figure 10: Predictive performance of the proposed model for the water level dataset considering predictions 1 to 5-days ahead under two different scenarios (unseen and observed locations). The dots represents the posterior mean and while the vertical bars represent the 95% prediction interval.
@@
Table 3: RMSE of OOS evaluation for prediction in 1 to 5-days ahead according to data availability (observed and unseen).
Date Observed location Unseen location
29/03/2022 (1-day ahead) 0.035 0.135
30/03/2022 (2-day ahead) 0.073 0.148
30/03/2022 (3-day ahead) 0.101 0.159
01/04/2022 (4-day ahead) 0.131 0.174
02/04/2022 (5-day ahead) 0.153 0.188
@@
Figure 11 depicts the predicted water level - after back-transformation to the original scale - for 20 unobserved locations. At each observed time point, an interpolation was performed to estimate water level at these unseen spatial points. For dates extending beyond the training dataset (29/03/2022 to 02/04/2022), the spatial latent field was first computed for the observed locations and subsequently interpolated in order to predict the water level at the unobserved locations, as outlined in Section 3.3. The limited information beyond 02/04/2022 is reflected in wider prediction intervals, indicating greater uncertainty in predictions.
Overall, the model demonstrated capacity to predict water level at unobserved locations, including for future days. It successfully captured the general temporal patterns of water level variation across most unseen locations. However, for some unobserved locations (stations 18002, 18019, 23030, 25001 and 29002), the predictions were conservative and failed to capture the scale of the variability of water levels, particularly, struggling to accurately predict the peak events. Additionally, at some sites (station 14121 and 22009), the model overestimate the water level fluctuations. These inconsistencies suggest that while the model generally performs well for unobserved locations, further refinement may be necessary to accommodate localised hydrological behaviour and site-specific characteristics.
@@
Figure 11: Predicted water level - after back-transformation to the original scale - and its true values for 20 unobserved stations. The black dots represent the real values, while the blue line represents the posterior mean with the shaded area indicating the 95% prediction interval.
@@ 5 Model Comparison
Despite the model’s novelty in the hydrology field, we aim to compare its prediction performance and uncertainty quantification with other methodologies in the literature within two different setups: (1) a comparison with a competitor model for large spatiotemporal datasets and (2) a comparison with models focused on a single basin, a strategy commonly adopted by hydrologists. This approach treats each river system independently, without leveraging information from other basins. For this setup, we also assess the appropriateness of the specification of the proposed model by exploring an alternative formulation for the space-time effect.
@@ 5.2 Comparison with a single-basin model
In this section, we investigate the advantages of the proposed model by comparing it with models that focus on a single basin, as well adopted in the hydrology field (Varouchakis et al.,, 2022; Varouchakis and Hristopulos,, 2019; Kazemi et al.,, 2021; Ruybal et al.,, 2019). Single-basin models are convenient because they alleviate the computational bottleneck of inverting large covariance matrices while preserving the unique characteristics of the basin. However, they may limit the model’s ability to leverage information from multiple basins, which could enhance its performance. Our comparison aims to highlight the trade-off between computational feasibility and model robustness when scaling to larger, multi-basin datasets.
@@
GP 1: Autoregressive process into the dependent variable
Spatiotemporal GP model with mean parameter at current time depends on the response in the previous time.
$p(\boldsymbol{y}_{t}\mid\boldsymbol{y}_{t-1},\boldsymbol{X}_{t},\boldsymbol{X}_{t-1},\boldsymbol{\theta})\sim\mathcal{N}(\boldsymbol{y}_{t};\boldsymbol{\mu}_{t},\boldsymbol{\Sigma}+\tau^{2}\boldsymbol{I})$
@@
Spatiotemporal GP model with the temporal dependence is incorporated into the spatial process.
$p(\boldsymbol{y}_{t}\mid\boldsymbol{X}_{t},\boldsymbol{z}_{t},\boldsymbol{\theta})\sim\mathcal{N}(\boldsymbol{y}_{t};\boldsymbol{X}_{t}\boldsymbol{\beta}+\boldsymbol{w}_{t},\tau^{2}\boldsymbol{I})$
$\boldsymbol{w}_{t}=\rho\boldsymbol{w}_{t-1}+\boldsymbol{\xi}_{t},\quad\boldsymbol{\xi}_{t}\sim GP(\boldsymbol{0},\boldsymbol{\Sigma}).$ (12)